OpenStudy (anonymous):

evaluate expression for x=3 and y=2

OpenStudy (lieutenantgeneral):

whats the question?

OpenStudy (anonymous):

3x+5y+9

OpenStudy (lieutenantgeneral):

x = 5 n + 2 y = -3 (n + 1)

OpenStudy (anonymous):

evaluate this expression 3x+5y+9 x=3 and y=2

OpenStudy (lieutenantgeneral):

oh ok 19

OpenStudy (anonymous):

its wrong

OpenStudy (lieutenantgeneral):

ok srry

OpenStudy (anonymous):

do you know it

OpenStudy (anonymous):

it said to use pemdas